Aiko, Lani, and Charlie own the 3-Star Partnership, sharing profits and losses 20:50:30. During the current year, 3-Star has total gross income of $500,000 and total allowable deductions of $300,000. How should each of the following taxpayers account for 3-Star's results? Explain.

a. 3-Star Partnership

b. Aiko

c. Lani

d. Charlie
Answer #1

Answer:-

a) 3-star partnership = 500000-300000 =$ 200000

b)Aiko = 200000*(20/100) =$ 40000

c) Lani = 200000*(50/100) = $100000

d)Charlie= 200000*(30/100) = $60000